OpenStudy (anonymous):

v is in the interval (pi/2, pi), sin v=1/15. what is cos(2v)?

OpenStudy (shayaan_mustafa):

Hi gd0x :) look at your interval. It should be look like this. (pi, pi/2) or you have missed - sign (-pi/2, pi)

OpenStudy (anonymous):

ok you have \[\sin(v)=\frac{1}{15}\] and you need \[\cos(v)\] before you can find \[\cos(2v)\]

OpenStudy (anonymous):

adjacent side is \[\sqrt{15^2-1^2}=\sqrt{225-1}=\sqrt{224}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

since you are in quadrant 2, cosine is negative so \[\cos(v)=-\frac{224}{15}\] and now use \[\cos(2v)=2\cos^2(v)-1\]

OpenStudy (anonymous):

another typo damn should have been \[\cos(v)=-\frac{\sqrt{224}}{15}\]

OpenStudy (anonymous):

now use \[\cos(2v)=2\cos^2(v)-1\]

OpenStudy (anonymous):

can i give an exact answer using purely numbers?

OpenStudy (anonymous):

you should get \[\cos(2v)=2\times \frac{244}{225}-1\] whatever that is

OpenStudy (anonymous):

yes, it is the number i wrote above

OpenStudy (anonymous):

\[\frac{223}{225}\]

OpenStudy (anonymous):

yep, haha. that was it.

OpenStudy (anonymous):

thanks a lot.